OpenStudy (genny7):

Help? y^2 - 12y + 20

OpenStudy (anonymous):

Are you just combining like terms?

OpenStudy (genny7):

I think so?

OpenStudy (anonymous):

whats the actual question or does it just say solve?

OpenStudy (anonymous):

just distribute -12y so that you get a product of positive 20... you can then solve for y

OpenStudy (anonymous):

hope you got it ?

OpenStudy (genny7):

Well the actual question says to: factor completely. If the polynomial is not factorable, write prime

OpenStudy (anonymous):

it is factorable...

OpenStudy (genny7):

Ok so then -12 would factor down to -3

OpenStudy (nurali):

This is how you solve it. y^2-12y+20 Okay, you need two numbers one that has a product of 20 and one that gives you, when added, -12. This can be either two positives, two negatives, or a combination of the two. Now, solving it. We know that -10 and -2 give you a product of 20 and -10 and -2, added together, give -12. So, the answer would be (y-2)(y-10) Because y*y= y^2 y*-10= -10y -2*y=-2y and -2*-10=20 Combine like terms -10y and -2y you get -12y. Which leads us back to the original problem, y^2-12y+20. The answer (y-2)(y-10)

OpenStudy (genny7):

Well both of those numbers can be factored by 4

OpenStudy (genny7):

2+ 10 add up to 12

OpenStudy (alekos):

two numbers that add up to -12 and those same two numbers to have a product of 20

OpenStudy (alekos):

are you confused?

OpenStudy (genny7):

Ok i kind of got that part

OpenStudy (genny7):

So then would i be getting: (y-4) (y^2 + 2y + 10)

OpenStudy (alekos):

No that's not it another way of looking at it y^2 - 12y +20 = (y - a)(y-b) can you see that?

OpenStudy (alekos):

(y-a)(y-b) = y^2 - (a+b)y +ab

OpenStudy (alekos):

so we need to find a+b =12 and ab = 20

OpenStudy (alekos):

what would those two numbers be?

OpenStudy (genny7):

2+10=12 5 x4 =20

OpenStudy (mrnood):

@genny7 you were given the complete (correct) answer above by @Nurali (which is against the CoC) @genny7 It looks to me that you do not fully understand the process (or purpose) of factorising an expression. Maybe you should look back at your text and classwork and try to get the basic concepts clear. Factorising essentially produces 2 expressions that multiply together to give the SAME as your original expression.

OpenStudy (mrnood):

btw @alekos made the same mistake as I did it is -12 and 20 (NOT 12 and 20)

OpenStudy (genny7):

Oh now i see what @Nurali was saying...& well i have some notes but to be honest they confuse me

OpenStudy (mrnood):

This is used when oyu have an expression with x^2 in it. NOTE - this current question is NOT an equation - it cannot be 'solved' but it is an expression that can be re-arranged) SO we are looking for 2 expressions that multiply to give an expression with x^2 in it. The only way to get x^2 is x*x so the 2 expressions must be (x+something) * (x+somethiing else) OK so far?

OpenStudy (genny7):

Yeah i think im ok so far..i think i got it know

OpenStudy (alekos):

Genny. You were right with the first line of your answer 10+2 = 12, a + b = 12 then we go 10 x 2 = 20, ab = 20 so a = 10 and b = 2 putting these in the equation i wrote before and we have (y-a)(y-b) = (y-10)(y-2) = y^2 - 12y + 20

OpenStudy (alekos):

........or expanding (y-10)(y-2) = y^2 - (10+2)y + 10*2 = y^2 -12y +20 do you see what's happening?

OpenStudy (genny7):

Yes i do see whats going on

OpenStudy (alekos):

ok, good

OpenStudy (genny7):

Yeah, thanks so much for your time and help :) @alekos

OpenStudy (alekos):

your welcome.